# DiffGAN: A Test Generation Approach for Differential Testing of Deep Neural Networks

### v1

@@ 7 Related Work
The above test generation approaches heavily depend on the availability and quality of seed data, limiting the process to mutated versions of existing inputs. This dependence significantly narrows the range of test scenarios, as it prevents the introduction of entirely new features (i.e, features that are not present in the initial testing dataset) into the testing process—a capability that DiffGAN notably possesses, enhancing its effectiveness in differential testing by generating diverse and new test input scenarios.
@@ 8 Conclusion
We presented DiffGAN, a novel black-box test generation framework that leverages GANs and NSGA-II for differential testing of deep neural networks. DiffGAN uses customized fitness functions in NSGA-II to guide the search process, generating diverse and valid triggering inputs that expose behavioral differences between models. To the best of our knowledge, this is the first work addressing black-box differential testing aimed at model selection.
We show that DiffGAN is capable of revealing behavioural differences across models by exploring large input spaces through GAN and NSGA-II’s diversity-driven fitness functions. Our empirical evaluation on eight pairs of image classification models trained on two widely-used datasets demonstrated that DiffGAN significantly outperforms the SOTA differential testing baseline, DRfuzz, by generating on average four times more triggering inputs. These inputs are not only more diverse but also more likely to be valid, allowing for a deeper analysis of the models’ behavior. Furthermore, we showed how these inputs can be applied in a machine learning-based voting mechanism, which dynamically selects the most accurate model based on input characteristics, resulting in improved classification accuracy compared to the baseline approach.
For future work, we aim to explore DiffGAN in dynamic environments to assess the side effects of updating a single ML component on the overall system performance. Rather than comparing pairs of DNN models, we will compare the original system with the updated version. This approach will provide valuable insights into maintaining system reliability and adaptability after updates, particularly in critical areas like autonomous driving, and healthcare.

### v2

@@ 5 Empirical Results
6 Discussions
Relevance of our Approach: We present in this paper DiffGAN, a test generation approach for the differential testing of DNN models for image classification. DiffGAN leverages NSGA-II, a multi-objective optimization algorithm, to effectively explore the latent space of a GAN and generate diverse and valid triggering images. Unlike traditional mutation-based approaches that rely on modifying seed inputs [8, 9, 10], DiffGAN directly explores the high-dimensional GAN latent space, allowing it to generate entirely new images with novel features that can trigger behavioral disagreements between the models under test. NSGA-II plays a crucial role by balancing two fitness functions: (1) maximizing divergence—ensuring, to generate inputs leading to significant differences in models outputs, and (2) maximizing diversity, thus reducing the generation of redundant images. Our ablation study shows that these two fitness functions are complementary and both essential for the generation of triggering images. By integrating GAN-based test inputs generation with NSGA-II-driven optimization, DiffGAN not only generates more triggering inputs than the SOTA approach but also ensures that these inputs are valid and more diverse.
Practical Applicability: The black-box nature of DiffGAN makes it agnostic to the models under test, as it does not require any assumptions on their architectural similarity. This flexibility allows DiffGAN to be applied across diverse practical scenarios where models with similar accuracy levels are developed independently or trained using different data and optimization techniques. The ability of DiffGAN to generate diverse and valid triggering inputs makes it a practical approach for evaluating and comparing DNN models with similar accuracy levels. DiffGAN is therefore particularly valuable for model selection, regression testing, and model compression, that are common application scenarios in ML development pipelines. In model selection, the triggering inputs generated by DiffGAN can help identify behavioral discrepancies enabling fine-grained analysis beyond accuracy metrics, ensuring the selection of the most adequate model. In regression testing, DiffGAN can be used to generate cases where an updated model underperforms compared to its predecessor. In model compression validation, DiffGAN can be used to understand when compressed models fare less well than their original versions and whether this is acceptable. Another important application of DiffGAN is ML-based model selection, a scenario informed by our industrial partner. In some autonomous vehicles, classifiers with similar accuracy levels are dynamically selected on the fly to classify sensor-captured images. Since different classifiers specialize in distinct regions of the feature space, selecting the most suitable one for each input improves reliability. Additionally, certain classifiers require more computational resources than others, making dynamic selection crucial for optimizing inference time and energy consumption, particularly in embedded systems with limited resources. By generating diverse inputs that expose model discrepancies, DiffGAN enhances the accuracy of model selection mechanisms since their training dataset is enriched by such inputs and more informative. The relevance of this application is further supported by existing literature on ML-based model selection [51, 52].
Generalisability: Although DiffGAN is primarily designed for image classification tasks, our approach can be extended to other domains such as image regression and segmentation. This extension can be achieved by adapting the divergence fitness function. For instance, in regression tasks, triggering inputs can be defined as those where the difference between the models’ outputs exceeds a predefined threshold. The divergence fitness function can then be tailored to capture a meaningful divergence metric, such as absolute or relative differences between regression model outputs, depending on the specific context. The primary objective would be to maximize the discrepancy between numerical outputs. For segmentation tasks, triggering inputs can be determined based on differences in predicted labels, variations in segmented shapes, or discrepancies in segmented regions. Once a triggering input is defined, the divergence fitness function can be customized to highlight these differences in the models’ outputs effectively. We should note that, for both regression and segmentation, the diversity-based fitness function will remain unchanged as long as the inputs are images. However, if the input type varies, the fitness function should be adapted accordingly to ensure an appropriate measure of diversity. We should also note that DiffGAN is not applicable to tasks involving multi-modal inputs or time-series data in its current form. This arises from the fact that our framework relies on GANs, which are specifically designed for image generation and manipulation. Supporting multi-modal or time-series tasks would require several modifications, such as integrating alternative generative architectures (e.g., Variational Autoencoders or Recurrent Neural Networks) and designing new strategies for defining and optimizing triggering inputs. While DiffGAN is designed for image classification, its black-box differential testing approach can be extended to LLMs by adapting input generation using trained text-based models, employing embedding similarity or log-likelihood divergence as a fitness function, and integrating language model discriminators for validity filtering. As future work, we aim to extend DiffGAN to support a broader range of ML tasks beyond image classification. This includes adapting the fitness function to better capture model discrepancies and the diversity of the generated inputs. We also aim to explore how DiffGAN can be adapted for other types of inputs, such as textual and audio inputs.
GAN Mode Collapse Mitigation: One of the key challenges of GANs is their propensity for mode collapse, which limits the diversity of the generated inputs. To mitigate this issue, our approach integrates multiple strategies during both the training phase and test input generation to improve the generation of diverse inputs. Specifically, during training, we ensured that GANs were properly optimized, achieving low FID and KID scores (Section 3.2). These metrics, widely recognized for evaluating the quality and diversity of generated samples, were used to confirm that the trained GANs effectively captured the diversity of the underlying training data distribution [33]. Low FID and KID scores indicate that the GAN not only avoids mode collapse but also generates realistic and diverse samples. Another factor we used to improve the diversity of the data was generating an augmented training dataset to train the GAN by applying different image transformations (Section 3.1). This data augmentation exposed the GAN to a more diversified training dataset, mitigating the risk of mode collapse. Furthermore, during the test case generation process, we explicitly incorporated diversity as a key fitness function. As detailed in Section 3.3.2, we clustered generated images based on their features and ensured that new images were distinct from previously generated ones. This approach guided the GAN to explore a broader range of outputs, reducing the likelihood of mode collapse and maintaining high diversity in the generated triggering inputs. Finally, our evaluation results demonstrated that our approach consistently outperformed the baseline in generating more diverse triggering inputs, thus validating the effectiveness of our diversity-enhancing strategies.
Efficiency. To evaluate the performance of DiffGAN on large image datasets, we analyzed how GAN training time evolves with increasing dataset size. Specifically, we created larger-scale datasets by applying diverse image transformation techniques to the MNIST dataset, expanding the dataset size from 60,000 to 100,000 images. We then trained multiple GANs on these datasets, systematically increasing the number of training samples to assess DiffGAN’s scalability. For each training run, we recorded training time and evaluated the quality of generated images using FID and KID scores.
Figure 6: Relationship between dataset size and training time of Vanilla GAN.
As shown in Figure 6, we found that DiffGAN scales linearly with increasing dataset size, with a training time ranging from 371 to 607 seconds. We note that this increase remains practically acceptable, as GAN training is a one-time process per dataset, and differential testing is neither frequent nor a real-time task. Thus, training costs do not significantly impact the feasibility of using DiffGAN for large-scale differential testing. Additionally, it is important to clarify that DiffGAN does not strictly require training a GAN from scratch, as it treats GANs (or any other generative model) as black-box components. This means that, for large-scale datasets such as ImageNet, DiffGAN can leverage pretrained GANs or generators that have already been trained on such datasets [74, 75], significantly reducing the computational burden while maintaining its ability to generate diverse and valid test cases. Existing GAN models trained on large datasets can be seamlessly integrated into DiffGAN without requiring additional training, making our approach even more scalable and applicable beyond the datasets tested in our current study.
@@ 8 Related Work
The above test generation approaches heavily rely on mutation operations applied on seed images, thus limiting the process to mutated versions of existing inputs. This dependence significantly narrows the range of test scenarios, as it prevents the introduction of entirely new features (i.e, features that are not present in the initial testing dataset) into the testing process. In contrast, DiffGAN enhances its differential testing effectiveness by generating diverse and new test input scenarios.
@@ 9 Conclusion
We presented DiffGAN, a novel black-box test generation framework that leverages GANs and NSGA-II for differential testing of deep neural networks. DiffGAN uses customized fitness functions in NSGA-II to guide the search process, generating diverse and valid triggering inputs that expose behavioral differences between models. To the best of our knowledge, this is the first work addressing black-box differential testing aimed at model selection. We show that DiffGAN is capable of revealing behavioural differences across models by exploring large input spaces through GAN and NSGA-II’s diversity-driven fitness functions. Our empirical evaluation on eight pairs of image classification models trained on two widely-used datasets demonstrated that DiffGAN significantly outperforms the SOTA differential testing baseline, DRfuzz, by generating on average four times more triggering inputs. These inputs are not only more diverse but also more likely to be valid, allowing for a deeper analysis of the models’ behavior. Furthermore, we showed how these inputs can be applied in a machine learning-based voting mechanism, which dynamically selects the most accurate model based on input characteristics, resulting in improved classification accuracy compared to the baseline approach.
For future work, we aim to enhance the generalizability of DiffGAN by extending it to support a wider range of models, machine learning tasks, and input types. We also aim to explore DiffGAN in dynamic environments to assess the side effects of updating a single ML component on the overall system performance. Rather than comparing pairs of DNN models, we will compare the original system with the updated version. This approach will provide valuable insights into maintaining system reliability and adaptability after updates, particularly in critical areas like autonomous driving, and healthcare.